OpenStudy (babycease):

Solve: negative 1 over 2 x + 1 = –x + 8 7 1 over 2 14 18 Please Help

OpenStudy (legomyego180):

\[\frac{ -1 }{ 2x+1 } = -x+8\] Is this correct?

OpenStudy (legomyego180):

What do you think we should do first?

OpenStudy (babycease):

Im not sure really im new to this

OpenStudy (legomyego180):

The end goal here is to figure out what number you can plug into x to make this equation "true". If you get stuck on the test, and its multiple choice, a smart thing to do is just plug in the answers and see which answer makes the equation true. It's not the right way to do it but it will get you by. However, I'll try to explain this to you...

OpenStudy (babycease):

Okay

OpenStudy (legomyego180):

First thing to do is remember order of operations. Lets get rid of that fraction...it looks like division so to get rid of it we have to do the opposite of it to the other side. In other words we need to multiply EVERYTHING on the right hand side by (2x+1). \[-1 = (-x+8)(2x+1)\] After we've gotten rid of the fraction you need to FOIL these two expressions out. \[-1 = -2x^2-x+16x+8\] Next combine like terms and move the -1 over to the right side. \[-2x^2+15x+9=0\] ... Are you sure you wrote this question right?

OpenStudy (babycease):

yes positive

OpenStudy (legomyego180):

what class is this?

OpenStudy (babycease):

Nvm I guessed and got it

OpenStudy (legomyego180):

None of these answers could be correct.

OpenStudy (legomyego180):

oh well